Norio Manufacturing uses powdered plastics (PPS) to manufacture a high-pressure board used in a digital equipment product, Flex 10. Information concerning its operation in June is as follows:

Budgeted units of Flex 10 for June 6,400

Budgeted usage of PPS 57,600 pounds

Actual number of units of Flex 10 manufactured 5,400

PPS purchased 71,280 pounds

PPS used 53,000 pounds

Total actual cost of PPS used $ 321,180

Direct materials usage variance $ 35,640 unfavorable

Assume that Norio does not maintain an inventory of materials, so that the amount of materials used is equal to the amount of materials purchased. The direct materials purchase-price variance (rounded to the nearest dollar) is:

Multiple Choice

  • $71,508 favorable.

  • $80,978 favorable.

  • $145,411 favorable.

  • $88,708 favorable.

  • $103,680 favorable.
Answer #1

Answer:

Option 3rd is correct.i.e $145,411 favorable.

standard cost per pound= 53000- {5400(57600/6400)}= 4400
35640/4400= 8.1
actual cost per pound = 321,180/53000= 6.06
direct materials purchase-price variance = pounds of pps purchased* (standard cost per pound - actual cost per pound)
71280*(8.1 - 6.06)= 145,411 favourable